SpaceRef

SpaceRef


MRPO Daily Status Report 3-18-03

Status Report From: Microgravity Research Program Office
Posted: Tuesday, March 18, 2003


GMT Date: 077

Calendar Date: 03/18/03

Increment: 6

Stage: 11A

Week 16

Payloads currently on-board ISS:
BSTC, CGBA, CSLM-2, InSPACE, MAMS, MSG (facility), PFMI, SAMS, STES-10, SUBSA and ZCG.

Status:
MAMS, SAMS and STES-10 continue nominal, basically unattended, operations. BSTC, CGBA, CSLM-2, InSpace, PFMI, SUBSA and ZCG are inactive. MSG trouble-shooting is continuing.

EXPRESS RACKS 1 and 4 (ER1 & ER4) are powered and supporting payload operations. ER2, ER3 & ER5 are currently powered down.

Evaluation of the MSG circuit breaker problem continues, and troubleshooting procedures are being implemented. Thirteen sensor cables were disconnected from MSG on day 59, leaving five sensors connected that were verified to cause no problem with MSG powered. All of the disconnected cables have been reconnected one at a time, to hopefully isolate the troublesome circuit, but no anomalies have occurred. MSG troubleshooting continued today and included power cycling ESEM 4A outlet 2 (MLC), ESEM 4B outlet 2, and ESEM 4A outlet 1 (Video drawer) sequentially. Today’s tests had no apparent anomalous results. Further trouble-shooting is contingent on the results of an ESA Control Board meeting this Thursday.

InSpace Coil assemblies may need to be inspected again prior to beginning test operations if it becomes feasible to install the payload in the MSG.

MAMS is operating nominally in the quasi-steady mode. HiRAP is currently inactive. MAMS data is available on the acceleration measurement web page.

PFMI hardware removal from the MSG is contingent upon success of MSG trouble-shooting.

All SAMS sensors in EXPRESS RACKS are operating nominally. The SAMS sensor in the MSG is currently turned-off. SAMS data is also available on the acceleration measurement web page.

STES-10 health and status data is nominal. STES-10 is configured with diffusion-controlled chambers that require no crew operation activities.

Near-term Outlook:

Week 16 (3/17-3/23) – Planning for this week was based on approximately 22 hours of crew time for NASA payload operations. The crew spent additional time today completing the replacement of a thermal control loop pump that failed Sunday. The crew will continue preparations for the recently defined EVA now being planned for week 19. Further MSG trouble-shooting is contingent on the results of an ESA Control Board meeting this Thursday.

Week 17 (3/24-3/30) - Preliminary planning for this week is being based on approximately 2 hours of crew time for NASA payload operations. The crew will continue preparations for the recently defined EVA now being planned for week 19. Time can be allocated for MSG trouble-shooting if it is needed.

M. Vlasse, PhD, DSc
PSR Increment Scientist


ACRONYMS:

BCSS Biotechnology Cell Science Stowage
BSTC Biotechnology Specimen Temperature Controller (General status information on this payload will include coverage of the following items of support equipment: BTR, BCSS and GSM)
BTR Biotechnology Refrigerator
CGBA Commercial Generic Bioprocessing Apparatus
ESEM Exchangeable Standard Electronic Module
ER1 EXPRESS RACK 1
ER2 EXPRESS RACK 2
ER3 EXPRESS RACK 3
ER4 EXPRESS RACK 4
ER5 EXPRESS RACK 5
GSM Gas Supply Module
InSPACE Investigating the Structure of Paramagnetic Aggregates
MAMS Microgravity Acceleration Measurement System
MSG Microgravity Science Glovebox
OCR Operations Change Request
PAO Public Affairs Office
PFMI Pore Formation and Mobility Investigation
PGB Plant Generic Bioprocessing Apparatus
PSR NASA Code U Physical Sciences Research Division
RIC Rack Interface Controller
SAMS Space Acceleration Measurement System II
SPD NASA Code U Space Products Development Division
STES Protein Crystal Growth - Single Thermal Enclosure System
SUBSA Solidification Using a Baffle in Sealed Ampoules
ZCG Zeolite Crystal Growth

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.